OpenStudy (anonymous):

Find the cosine of angle A. 0.88 BC=8 AB=17 AC=15

OpenStudy (anonymous):



OpenStudy (shubhamsrg):

in right angled triangle,,whats cos of an angle mam ?

OpenStudy (apoorvk):

cosA= base/hypotenuse = AC/AB

OpenStudy (apoorvk):

(if it's a right angled triangle)

OpenStudy (anonymous):

\(\ \Huge cos=\frac{adj}{hyp}=\frac{AC}{AB}=\frac{15}{17} \).

OpenStudy (anonymous):

im asking if that the way you write these answer with a zero at first 0.88

OpenStudy (shubhamsrg):

tell me this,,which is greater ,, 0.88 , 00.88 , .88 ?

OpenStudy (anonymous):

dont matter

OpenStudy (shubhamsrg):

huh ?

OpenStudy (anonymous):

i mean they are the same

OpenStudy (shubhamsrg):

yes..exactly.. so it dosent matter if you write 0.88 or .88 or anything of that sort..

OpenStudy (anonymous):

ok thx